### v1

# Rolling the DICE on Idiomaticity: How LLMs Fail to Grasp Context

Abstract
Human processing of idioms relies on understanding the contextual sentences in which idioms occur, as well as language-intrinsic features such as frequency and speaker-intrinsic factors like familiarity. While LLMs have shown high performance on idiomaticity detection tasks, this success may be attributed to reasoning shortcuts in existing datasets. To this end, we construct a novel, controlled contrastive dataset designed to test whether LLMs can effectively use context to disambiguate idiomatic meaning. Additionally, we explore how collocational frequency and sentence probability influence model performance. Our findings reveal that LLMs often fail to resolve idiomaticity when it is required to attend to the surrounding context, and that models perform better on sentences that have higher likelihood. The collocational frequency of expressions also impacts performance. We make our code and dataset publicly available. [^1] https://github.com/mi-m1/dice
Rolling the DICE on Idiomaticity: How LLMs Fail to Grasp Context
Maggie Mi1 Aline Villavicencio1,2 Nafise Sadat Moosavi1 1University of Sheffield 2University of Exeter zmi1@sheffield.ac.uk a.villavicencio@exeter.ac.uk n.s.moosavi@sheffield.ac.uk
1 Introduction
Idiomatic expressions (IEs) are strange birds whose meaning may not be straightforwardly related to the meaning of the component words in isolation. For example, proficient English speakers understand “spill the beans” not as causing legumes to fall, but as disclosing a secret. Estimates suggest that there are 25,000 fixed expressions in English alone (Wèinreich, 1969), and a similar estimate is quoted for French (Gross, 1982). Notably, this figure is comparable to the order of magnitude of individual words in the lexicon (Jackendoff, 1997). This suggests that idioms are not mere linguistic curiosities but fundamental components of language.
Additionally, some of these are ambiguous, “potentially idiomatic expressions" (PIEs) that can be interpreted either non-compositionally (figuratively or idiomatically) or compositionally (literally), depending on the context in which they appear. Accurately identifying the meaning of a PIE often depends on its context and is essential for numerous downstream applications, such as machine translation (Dankers et al., 2022; Barreiro et al., 2013; Salton et al., 2014; Fadaee et al., 2018), sentiment analysis (Williams et al., 2015; Liu et al., 2017), and automatic spelling correction Horbach et al. (2016). Beyond these applications, it is also crucial to grasp the underlying meaning of the text.
Existing datasets that feature expressions with both literal and idiomatic usages often fail to rigorously examine the role of context in disambiguating meaning. This shortcoming arises because the literal meanings in such datasets typically result from syntactic modifications or semantic shifts, which inherently disrupt the idiomaticity of the expression. For example, passivization (2) or modification to the expression (3) often strip the expression of its idiomatic meaning, as shown by the idiomatic usage in (1). Consequently, this allows models to exploit surface-level differences, such as changes in grammatical structure or added modifiers, as shortcuts for determining literalness, rather than relying on their understanding of idiomaticity in context as the task intends. This undermines the evaluation’s effectiveness by encouraging models to use superficial cues instead of deeper contextual comprehension.
[…]
In this paper, we address this gap by introducing a new benchmark designed to evaluate the of ability of LLMs to interpret idiomatic expressions in context. We focus on idioms, as these figurative expressions may serve as key indicators of a model’s linguistic understanding. Given the possibility of a predominant sense, our dataset presents idiomatic expressions, balancing both literal and figurative contexts, challenging models to rely on contextual cues for accurate interpretation. This contrastive evaluation forces models to distinguish between meanings based solely on context, preventing them from relying on memorized idioms. We hypothesize that if models do not depend on memorization, they will perform equally well in both senses of a PIE.
Moreover, in human language processing, frequency (Swinney and Cutler, 1979) and familiarity (Nippold and Rudzinski, 1993), alongside context (Estill and Kemper, 1982; Gibbs Jr et al., 1989; Gibbs and Nayak, 1989), are known to influence idiom comprehension and speed of processing (Tabossi et al., 2009). This may similarly apply to models. In addition to context, we analyze the impact of a language-intrinsic feature (expression frequency) and a model-intrinsic feature (likelihood) on model performance.
Contributions
(1) We present a comprehensive and robust evaluation dataset, containing PIEs no syntactical changes appearing in both figurative and literal settings. (2) Through fine-grained evaluations, we find that models struggle to use context for idiomaticity processing. (3) Based on frequency estimates, we find frequency is not a “free lunch” and that, whilst highly frequent idioms maybe be more likely to be disambiguated correctly, there is a trade-off between performance on the two idiomaticity settings (literal/figurative). (4) For models that learned idiomaticity sufficiently well, we observe that there is a strong relationship between the likelihood of the contextual sentence and performance on idiomaticity detection.
2 Related Works
Existing Datasets
The biggest dataset for idiomatic sense disambiguation is MAGPIE (Haagsma et al., 2020). MAGPIE contains a total of 56,622 PIE instances, across 1,756 idioms. However, a large amount of deviation of the form of an expression was allowed when MAGPIE was curated. As a result, most of the literal uses of PIEs involve modifications to the form of the expression (similar to the example in Section 1). Other large datasets targeting various types of IEs have been released: The VNC-Tokens dataset focusing on verb-noun combinations Cook et al. (2008), IDIX on verb-noun phrase or verb-prepositional phrase expressions (Sporleder et al., 2010). SemEval-2013 has unrestricted expressions (Korkontzelos et al., 2013), and AStitchInLanguageModels contains noun compounds (Tayyar Madabushi et al., 2021).
To address these problems: (1) we propose a novel evaluation set (DICE) where we strictly control the form of idiomatic expressions. This eliminates the possibility that models rely on grammatical variations for idiomaticity disambiguation. By maintaining consistent forms across contexts, our dataset ensures that the challenge lies in understanding contextual nuances, thereby providing a more accurate assessment of a model’s idiomatic comprehension. (2) Visibly, these datasets often only contain expressions of a singular type. As a result, we address this lack of coverage by including expressions that are both phrasal expressions and noun compounds. Moreover, we believe it is crucial to maintain the same form of expression in both literal and figurative contexts, so we aim to construct a more controlled evaluation framework.
Idiomaticity Detection
The task of idiomaticity sense disambiguation (ISD), or, idiomaticity detection involves evaluating whether an expression is used literally or figuratively in a sentence (Liu and Hwa, 2018; Salehi et al., 2014; Senaldi et al., 2016; Gharbieh et al., 2016). LLMs seem to achieve high performance in idiomaticity detection tasks (Phelps et al., 2024; Zeng and Bhat, 2021). This task can be framed as a binary classification, where models must determine whether the usage of a PIE is literal or idiomatic. However, their ability to accurately model context relevant for idiomatic understanding is unclear Garcia et al. (2021b) and has yet to be comprehensively determined. Given the crucial role of context in interpreting these expressions, it is important to evaluate to what extent models use context to make this distinction. However, since these models are trained on extensive corpora that likely include idiomatic expressions, it is unclear whether they are memorizing the surface forms or genuinely comprehending the context to detect the idiomatic usage.
Contrastive Evaluation
Contrastive evaluation often takes the form of minimal pairs evaluation, where a single perturbation such as a change in a word or condition is systematically introduced into otherwise identical conditions Prasad et al. (2019); Garcia et al. (2021a). This method has been noted for its advantage in identifying specific weaknesses in model understanding and robustness (Linzen et al., 2016; Sennrich, 2017; Robertson, 2019).
Our work fits within this category of contrastive evaluation, with a specific focus on context comprehension in idiomaticity detection. By presenting PIEs in both their figurative and literal forms, our dataset forces models to make use of subtle contextual cues to differentiate between meanings. Additionally, by controlling for the form of the expressions, we ensure they remain identical across contexts. Thus, the challenge stems from interpreting the surrounding context, rather than recognizing variations in form.
Memorisation and Context
Pretrained language models tend to handle IEs mainly through memorization of stored expressions and token distributions rather than reasoning about the meaning in context. This reliance on memorization becomes evident when models face novel noun compounds. For instance, Li et al. (2022) found that while GPT-3’s interpretations of novel compounds closely matched human responses, its performance faltered when tasked with interpreting nonsensical strings. Unlike humans, who can use surrounding context to deduce meaning, GPT-3 failed to do so, indicating its heavy dependence on pre-learned token distributions rather than contextual clues. Similarly, Coil and Shwartz (2023) explored noun compound interpretation using GPT-3 and found that, although the model performs well with common noun compounds, its performance drops significantly with novel compounds. This suggests that the model relies on pre-existing knowledge for familiar compounds, but struggles when required to reason about unseen combinations. Their analysis highlights the tension between reasoning and parroting in large models, making a case for the limits of model comprehension on noun compound tasks.
Further illustrating the limitations of context utilization, Cheng and Bhat (2024) show that pretrained LLMs often perform better in idiomatic expression reasoning when context is removed entirely. This counterintuitive result suggests that context can sometimes confuse models rather than help them. Moreover, Sun et al. (2021) found that LLMs rely on context only when the answer can be directly retrieved from it. These results suggest that reasoning performance almost always increases with context removal.
3 DICE: Dataset for Idiomatic Contrastive Evaluation
We introduce ’DICE’, a novel dataset for controlled evaluation of potentially idiomatic expressions balanced for sense. An overview of the curation process adopted for DICE is presented in Figure 1.
Figure 1: A visual illustration of our dataset curation process. We obtain a list of idioms using existing datasets, which GPT-4 is then prompted to generate sentences where the idiom is used literally. Expert then annotate the sentences to select the literal annotation in the literal subset, and that the figurative subset consists of the examples from existing datasets.
Counts Examples and Remarks Number of Sentences (Literal) 1033 Carpenters recommend not to sand against the grain as it can damage the wood. Number of Sentences (Figurative) 1033 Out of duty she had caved in, but it still went against the grain. (MAGPIE) Total no. of sentences 2066 - Number of Unique Idioms 402 - Total Number of Expressions 402 103 noun compounds + 299 phrasal expressions Average length of sentences (literal) 15.4 words - Average length of sentences (figurative) 28.1 words -
Table 1: This table presents the characteristics of the subset of data used in our experiments and analysis. We have also gathered additional data, and a summary of that subset is available in Appendix B.
Expression Selection
To build our dataset, we compiled a comprehensive list of idiomatic expressions, focusing on both phrasal idioms and noun compound idioms. For phrasal idioms, we identified overlapping expressions from two established resources: MAGPIE (Haagsma et al., 2020) and SLIDE (Jochim et al., 2018). For noun compound idioms, we selected idiomatic expressions that were common to the NCTTI dataset (Garcia et al., 2021b) and AStitchInLanguageModels (Tayyar Madabushi et al., 2021). We specifically focused on idiomatic noun compounds, as the meaning of these expressions cannot be directly derived from the meanings of their individual words.
We excluded compositional and partially compositional compounds, as these tend to have dominant literal meanings that are difficult to override in context (e.g., “skin tone” or “noble gas”). Such expressions do not present the same interpretive challenge for models, as their meanings are more directly tied to their components. By focusing solely on non-compositional idioms, we ensure that the dataset tests the model’s ability to interpret figurative language based on context rather than relying on straightforward lexical composition. This selection process resulted in a total of 783 unique idiomatic expressions, comprising 680 phrasal idioms and 103 idiomatic noun compounds.
Sentence Generation
We used GPT-4 (OpenAI et al., 2024) to generate sentences where a given idiom appears in a literal context, intentionally suppressing its figurative interpretation. The specific prompting setup we used for sentence generation is provided in Appendix C. To determine the best model for sentence generation, we piloted the process using three versions: GPT-4o, GPT-4, and GPT-3.5. The models were prompted to produce three different sentences, where the form of the idiom must be kept the same. We tested out the annotation process (detailed below) on a set of 20 idioms, and the Cohen’s kappa coefficient of inter-rater reliability for this pilot stage was 0.987. Overall, we found GPT-4 to perform the best at generating sentences where the figurative interpretation is suppressed. Our preference for GPT-4 aligns with the findings of Phelps et al. (2024), which demonstrate that off-the-shelf GPT-4 possesses relatively stronger idiomaticity knowledge as it performed consistently well across idiomaticity detection tasks compared to other off-the-shelf LLMs. In total, we obtained 2,349 sentences from GPT-4.
Expert Annotations
We recruited four experts with at least three years of university-level experience in Linguistics, compensated at a rate of £15/hour. The annotators were tasked with reviewing each sentence and deciding whether to accept it unconditionally, skip it, or reject it if the idiom’s figurative meaning could not be fully suppressed. In cases of rejection, annotators provided reasons such as ambiguity, figurative interpretation, change of form, or other issues. We provide information about this subset in Appendix B and also the briefing the annotators received in Appendix D. If an expression was skipped, a second annotator reviewed it to confirm if it should be discarded. Examples of sentences for each category are presented in Table 3. The reliability of the annotation process was high, as reflected by a Cohen’s kappa coefficient of 0.95, indicating strong inter-rater agreement.
The figurative counterparts of these sentences were sourced from MAGPIE and AStitchInLanguageModels. We ensure that the same number of variants is matched between the figurative and literal settings. In other words, if we have three sentences containing “all hell broke loose” in literal contexts, we would extract an equal number of sentences containing the idiom from the figurative datasets. This ensures that the dataset is balanced with regards to the idiomatic and literal interpretation of each expression.
3.1 Dataset Statistics
In total, our contrastive evaluation dataset consists of 2,066 sentences, featuring 402 expressions. A summary of the statistics of our dataset is presented in Table 1. In this paper we only use the subset whose literal leaning was confirmed by the annotators, however, we also release the excluded sentences to enable further analysis. However, we release the complete dataset, with an additional subset that can serve as a resource for additional analyses and for creating even more challenging datasets.
4 How Well Do LLMs Use Context for Idiomaticity?
Using DICE, we evaluated the ability of various language models to differentiate between literal and figurative uses of idioms. Replicating the Idiomaticity Sense Disambiguation (ISD) task, we prompted each model with a sentence containing an idiom and asked it to classify the idiom as either “literal” or “figurative” based on its usage in context.
4.1 Experimental Setup
Models
We evalute 13 models on the task of idiomatic sense disambiguation. These models are: GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o, GPT-3.5-Turbo Brown et al. (2020), FLAN-T5 models in the XXL, XL, Large, Small sizes (Chung et al., 2023), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Llama 3 models (incl. Llama 3.1 (405B, 8B, 70B) (Dubey et al., 2024). Additionally, we evaluated GPT-4 (OpenAI et al., 2024), which was used to generate the sentences. The prompts and computational resources used for the experiments are reported in the Appendix E.
To ensure a comprehensive and dependable assessment, each model was given three distinct yet semantically equivalent prompt versions. We assessed the models under both zero-shot and few-shot conditions.
Evaluation
To thoroughly evaluate the models’ performance, we employed three distinct evaluation settings:
  • $\bullet$
    Accuracy: This setting includes two sub-evaluations: (1) Figurative Accuracy: We computed the accuracy of each model in correctly identifying the figurative uses of expressions within the figurative subset. (2) Literal Accuracy: We assessed the accuracy of the models in correctly identifying the literal uses of expressions within the literal subset. These evaluations measure the models’ ability to recognize idiomatic and literal meanings based on context.
  • $\bullet$
    Lenient Consistency: This metric rewards the model for consistently classifying all instances of an expression as either literal or figurative within a specific type. For example, if the model correctly classifies all literal variations of a given expression as “literal”, it earns a point for that expression. Similarly, the model is rewarded if it correctly identifies all figurative instances of an expression as “idiomatic”.
[…]
    $\frac{\sum_{x}\mathbf{1}\left(\forall i,\text{Pred}(x_{i}^{Lit})=\text{Lit})+\mathbf{1}(\forall i,\text{Pred}(x_{i}^{Fig})=\text{Fig}\right)}{2*\text{Number of unique expressions}}$ (1)
    where $x$ is a unique expression in the DICE, and $\mathbf{1}(\cdot)$ is the indicator function.
  • $\bullet$
    Strict Consistency: This is the most stringent evaluation. The model had to correctly identify all variations of an expression in both figurative and literal contexts to be rewarded. This setting assumes that a truly understanding model should correctly classify an idiom regardless of its context.
[…]
    $\frac{\sum_{x\in\mathcal{X}}\mathbf{1}\left(\forall i,\text{Prediction}(x_{i})=\text{True Label}(x_{i})\right)}{\text{Number of unique expressions}}$ (2)
By employing these evaluation settings, we aim to provide a comprehensive assessment of the models’ capabilities in understanding and differentiating idiomatic expressions. This approach helps us determine whether the models rely on contextual understanding or memorized patterns to perform the task.
Model Accuracy Lenient Consistency Strict Consistency
Figurative Literal Overall Figurative Literal Overall Both Settings
GPT-4o 87.05 ± 3.62 87.30 ± 2.98 84.33 ± 4.44 69.49 ± 11.71 71.06 ± 6.68 70.32 ± 7.11 48.59 ± 9.75
GPT-3.5 Turbo 79.05 ± 5.01 70.02 ± 12.72 75.54 ± 7.81 82.59 ± 9.17 44.36 ± 22.28 63.47 ± 7.61 32.84 ± 15.81
Flan-T5-XXL (11B) 77.18 ± 1.40 74.91 ± 8.35 76.40 ± 4.49 63.93 ± 13.71 58.79 ± 23.16 61.36 ± 4.73 32.92 ± 6.80
Flan-T5-XL (3B) 70.48 ± 3.56 33.94 ± 26.91 59.65 ± 8.19 91.13 ± 6.97 13.02 ± 11.24 52.07 ± 3.58 9.95 ± 8.88
Flan-T5-Large (780M) 66.63 ± 0.10 3.45 ± 4.72 50.42 ± 0.53 97.68 ± 3.40 0.58 ± 0.80 49.13 ± 1.30 0.58 ± 0.80
Flan-T5-Small (80M) 0.51 ± 0.59 66.72 ± 0.07 50.13 ± 0.15 0.00 ± 0.00 100.00 ± 0.00 50.00 ± 0.00 0.00 ± 0.00
Llama 3.1 (405B) 88.63 ± 2.36 88.25 ± 3.93 88.45 ± 3.10 78.52 ± 5.61 80.02 ± 12.43 79.27 ± 3.46 60.36 ± 6.61
Llama 3 (70B) 87.72 ± 4.63 86.13 ± 7.10 87.00 ± 5.73 81.84 ± 4.00 72.64 ± 16.12 77.24 ± 7.45 57.55 ± 12.41
Llama 3 (8B) 79.27 ± 1.97 74.01 ± 2.79 76.91 ± 2.25 77.86 ± 5.18 48.76 ± 3.37 63.31 ± 1.43 33.83 ± 2.60
Llama 2 (70B) 76.28 ± 4.39 56.64 ± 17.13 69.62 ± 7.82 93.20 ± 4.75 24.54 ± 16.89 59.12 ± 5.78 21.81 ± 13.51
Llama 2 (13B) 68.99 ± 1.39 36.09 ± 3.85 58.26 ± 1.96 85.41 ± 3.56 8.37 ± 3.34 46.93 ± 2.30 5.64 ± 2.00
Llama 2 (7B) 55.51 ± 19.54 31.97 ± 24.25 51.34 ± 1.55 59.87 ± 46.26 18.08 ± 29.16 38.97 ± 8.59 1.66 ± 1.37
GPT-4 88.56 ± 2.03 88.63 ± 2.08 88.48 ± 2.18 79.02 ± 3.11 78.03 ± 4.60 78.52 ± 2.95 59.62 ± 4.67
GPT-4o 89.43 ± 1.23 90.15 ± 1.71 89.72 ± 1.45 74.63 ± 1.99 87.40 ± 5.81 81.01 ± 1.93 63.52 ± 3.15
GPT-3.5 Turbo 79.41 ± 4.19 72.69 ± 10.87 76.70 ± 6.54 78.44 ± 8.80 49.42 ± 18.96 63.93 ± 5.92 34.16 ± 12.19
Flan-T5-XXL (11B) 10.20 ± 15.69 67.90 ± 1.91 52.79 ± 4.34 1.58 ± 2.52 99.25 ± 1.29 50.41 ± 0.61 1.49 ± 2.37
Flan-T5-XL (3B) 0.64 ± 0.80 66.71 ± 0.11 50.13 ± 0.22 0.08 ± 0.14 99.83 ± 0.29 49.96 ± 0.19 0.08 ± 0.14
Flan-T5-Large (780M) 3.28 ± 3.64 66.27 ± 0.45 50.00 ± 0.00 0.66 ± 0.76 96.93 ± 3.73 48.80 ± 1.48 0.00 ± 0.00
Flan-T5-Small (80M) 45.23 ± 39.19 35.55 ± 33.55 53.03 ± 5.25 60.78 ± 53.37 37.31 ± 54.62 49.05 ± 1.65 2.40 ± 4.16
Llama 3.1 (405B) 89.57 ± 1.80 89.54 ± 2.54 89.53 ± 2.17 79.10 ± 3.26 82.01 ± 7.85 80.56 ± 2.56 63.27 ± 4.66
Llama 3 (70B) 87.75 ± 3.76 86.97 ± 5.64 87.27 ± 4.61 78.52 ± 3.59 75.62 ± 14.01 77.07 ± 6.00 57.55 ± 10.22
Llama 3 (8B) 80.32 ± 5.33 73.81 ± 11.40 77.59 ± 7.62 79.35 ± 1.08 48.01 ± 15.70 63.68 ± 7.34 34.91 ± 13.59
Llama 2 (70B) 70.40 ± 1.19 31.44 ± 6.18 58.65 ± 2.28 96.52 ± 0.66 7.55 ± 2.75 52.03 ± 1.50 6.72 ± 2.63
Llama 2 (13B) 70.64 ± 1.15 34.20 ± 6.92 59.36 ± 2.45 94.94 ± 0.52 9.54 ± 4.14 52.24 ± 1.83 8.29 ± 3.11
Llama 2 (7B) 70.26 ± 3.14 42.18 ± 26.31 61.21 ± 9.28 80.76 ± 15.43 20.73 ± 22.25 50.75 ± 3.46 11.69 ± 10.42
GPT-4 88.52 ± 1.49 88.95 ± 2.09 88.42 ± 1.73 78.44 ± 0.76 77.94 ± 5.84 78.19 ± 2.63 58.87 ± 4.86
Table 2: Results $\pm$ 1 std (over 3 different sets of prompts) in zeroshot (top) and oneshot (bottom) conditions. Bold values represent the best performance for overall metric (incl.Strict Consistency) across models. GPT-4 is separated from the rest, as this model generated the evaluation sentences.
4.2 Results
Table 2 presents the results of model performances on our evaluation set. In it important to note that, GPT-4 was used to generate the sentences for DICE, so we do not consider its performance to be revealing.
Accuracy Comparison
With the exception of the Flan models (small to XL), which generally struggle with recognizing the idiomatic use of expressions, the other models demonstrate better performance in identifying figurative rather than literal usage. Notably, the GPT models and the LLama 3.1 (405B) model exhibit comparable results across both figurative and literal subsets, standing out as exceptions to this trend. A similar pattern is observed in both zero-shot and few-shot settings, although the one-shot setting consistently yields higher overall scores.
FLAN-XXL shows reasonable zero-shot performance but unexpectedly drops in few-shot settings, unlike GPT and LLama, which improve with more examples. We hypothesize that this decline is due to Flan models being heavily instruction-tuned and may require more specific prompt structuring in few-shot scenarios, such as clearer task-example separation or tailored formatting, compared to more advanced models like GPT and LLaMA that handle generalized prompting better.
Lenient Consistency
The general trend for this evaluation metric aligns with our previous observations: models show a preference for figurative interpretations when encountering an idiom, as there is a higher proportion of idioms that the models can consistently predict to be figurative across all contextual sentences than in the literal setting. As expected, all models achieve lower scores on both subsets when evaluated based on consistency, where the model must correctly classify all variations of the same expression in each sense to be rewarded. We observe the largest aggregate drop for GPT-4o when scores are evaluated using this metric. This decline indicates that GPT-4o’s high performance (evidenced by a overall accuracy of 84.33 ± 4.44 in zershot evaluations) stems from its consistency across a broad range of idioms. However, the model lacks a deep understanding of these idioms, making it susceptible to variations. This is illustrated by an overall Lenient Consistency score of 70.32 ± 7.1, showing that the model can only accurately interpret a subset of idioms consistently across different texts. Llama 3.1 is the model with the least performance difference across the two subsets, indicating a more balanced understanding of both figurative and literal contexts.
Strict Consistency
Strict Consistency, as previously defined, requires models to correctly classify all the figurative and literal uses of an expression to be rewarded. The results from this evaluation are striking: only three models—Llama 3.1, Llama 3, and GPT4o—achieve an accuracy above 40%, with 60.36%, 57.55%, and 48.59% respectively. This indicates that while state-of-the-art models may show high performance on existing idiomaticity benchmarks, they perform very poorly when a more systematic approach is used to evaluate their understanding. Even GPT-4, which served as the annotator model, can consistently classify only 59.62% of the expressions correctly in both literal and figurative contexts. This highlights a significant gap in the current models’ ability to truly understand idiomatic expressions.
5 Impact of Frequency and Sentence Likelihood on Model Performance
Refer to captionRefer to captionRefer to caption
Figure 2: Frequency results of GPT-3.5 Turbo, Llama 2 (70B), Flan-T5 XL (left to right).
This section examines how idiom frequency and sentence likelihood impact LLM performance. In human processing, idiom frequency and familiarity, alongside context, influence comprehension. Similarly, in LLMs, expression frequency and sentence probability may shape idiom detection. Exploring these factors helps clarify how language-intrinsic and model-intrinsic features affect performance.
5.1 Frequency Estimation
We use the English Web Corpus (enTenTen) (Jakubíček et al., 2013) to approximate the frequency of idioms in our dataset. This was for two key reasons: (1) it is parsed and tagged which allows us to query all morphological forms of the expressions using lemmas, ensuring comprehensive frequency counts, and (2) its large scale—52 billion words across diverse genres offers a robust generalization of natural language usage. While enTenTen is not identical to the pretraining datasets of LLMs, it provides a reasonable proxy for estimating expression frequency, as high-frequency terms in enTenTen are likely to appear frequently in similar web-based pretraining corpora. This analysis helps us understand whether model performance is affected by how often they may have encountered each expression during training.
Figure 3: Frequency distribution of idioms in DICE below 200,000 counts.
As shown in Figure 3, the frequency of idioms in DICE varies, with the majority of expressions occurring fewer than 200,000 times. The highest concentration of idioms falls within the 0 to 100,000 range. To focus on the most relevant portion of the dataset, we limit our analysis to this range to avoid skewed results.
Refer to captionRefer to caption
Figure 4: Likelihood results from Llama 3 (8B) and Flan-T5 XXL (left to right).
5.2 Likelihood Scores
LLMs make predictions using likelihood estimates learned during training, which are designed to maximize the probability of their extensive training data. Recent studies have shown that models tend to perform better, or assign higher evaluation scores, to sentences they deem more likely, regardless of the specific target task or evaluation criteria (Ohi et al., 2024; McCoy et al., 2024). This is likely due to sentences with higher likelihood being encountered more frequently during training. This can affect the model’s performance, as it is more familiar with some sentences than others. This in turn highlights the limitations of the models when dealing with less common contexts.
5.3 Results
To explore the impact of these two features on model performance, we fit a linear regression model, with idiom frequency/likelihood as the independent variable and the model’s accuracy as the dependent variable. This approach allows us to estimate how these two variations influence the ability of language models to distinguish between literal and idiomatic uses.
We present the frequency results for GPT-3.5 Turbo, Llama 2 (70B) and Flan-T5 XL in Figure 2. For the likelihood analysis, we present the results for Llama 3 (8B) and Flan-T5 XXL in Figure 4. Results for the other models can be found in Appendix F. These models were chosen as they are representative of the general patterns observed across all the models evaluated.
Frequency is not a free lunch!
In general, we observed a consistent pattern that accuracy correlates with expression frequency, suggesting that while models may have encountered certain expressions frequently during pretraining, they may still struggle to properly interpret their idiomaticity in new contexts. Notably, this trend was observed in 9 out of the 13 models we analyzed. We also observe that, as the frequency of an expression increases, the models tend to perform better at identifying its literal occurrences, but their accuracy in recognizing idiomatic uses declines. This could potentially be due to the literal meaning being the dominant occurrence of high-frequency expressions in the pretraining data; however, this hypothesis requires further investigation.
Sentence Likelihood
In the case of the Llama models, performance correlates positively with sentence probability. In both settings, the model perform better on sentences on which it has a higher likelihood. This is particularly the case on the literal subsets, as indicated by Llama 2 (13B) and Llama 3 (8b). For the Flan-T5 models, we see a negative or negligible correlation between frequency/likelihood and performance. As seen in §4.2, Small and Large does not appear to have effectively utilized the context to learn meanings as successfully as the other models. The counter-intuitive pattern observed in XL and XXL models could be due to models being over-confident in their wrong prediction. This leads to situations where model would assign high probabilities to idiomatic sentences but performs poorly on the idiomaticity detection task.
6 Conclusion
In this work, we contribute to idiomaticity detection in NLP with several key findings. First, we introduce DICE, a challenging dataset for context-dependent idiom detection, where distinguishing figurative from literal meanings relies heavily on understanding context. Second, we propose an evaluation framework that measures both overall accuracy and strict consistency, requiring models to correctly identify all figurative and literal instances of an expression across different contexts. Third, we show that current LLMs struggle to use context effectively, highlighting the need for models that better capture contextual nuances. Lastly, we find that while expression frequency aids recall, it is insufficient for idiomaticity detection without context, in both human cognition and model performance. We also observe a positive correlation between a model’s likelihood score and its performance on DICE, suggesting training exposure influences results, emphasizing the importance of context comprehension in LLMs.
7 Limitations
One of the limitations of our work is that some idiomatic expressions are noticeably more reliant on the context than others. This means that there were cases, where we could not provide a literal counterpart to the figurative interpretation. For example, the expression "set eyes on" has such a dominant meaning of "to see", that the annotators believed to be impossible to override. In these cases, we would discard the expression. As a result, our dataset only contains a selected sample of idioms, and we acknowledge that this idea of contrastive evaluation cannot necessarily be applied to all idioms in a language.
Another of the limitations of our work is that we only consider English idioms. We would like to have extended this work to other languages, however, this relies on the existence of idiomaticity datasets in the target languages. Moreover, the idea of making idioms literal might not be translatable to other languages, where the expression takes rigid and fixed forms.
8 Ethical Considerations
We adhere to ethical practices of data collection. All participants were required to sign a consent form and informed that they could withdraw from participation at any time without facing any consequences. Our collection procedures and processes are monitored and reviewed by a University-wide ethics committee. The committee members are unrelated and detached from this work.
Acknowledgements
We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1], UK EPSRC grantEP/T02450X/1, the Turing Institute Fellowship and the UniDive COST Action. Additional thanks to our annotators for their hard work. A further acknowledgement to Lily Zeng, Thomas Pickard and members of the MWE Group for their helpful discussions.
Appendix A Examples of Expert Annotations
We provid an example of expert annotations in Table 3.
Idiom Definition of the Figurative Meaning Sentence Accept Reject Reason (if reject)
smoking gun "a piece of incontrovertible evidence" The detective found a smoking gun at the crime scene. N Y Ambiguous
guilt trip "to make someone feel guilty" After breaking her mother’s vase, Sarah’s sister put her on a guilt trip for weeks. N Y Doesn’t make sense
turn a blind eye "pretend not to notice" Despite the obvious safety hazards, the supervisor chose to turn a blind eye. N Y Figurative
down the wire "a situation whose outcome is not decided until the very last minute" The electrician was careful not to cut down to the wire while he was working. N Y Form changed
set eyes on "see" As soon as she set eyes on the beach, she was overwhelmed by its serene beauty. N Y Skip
blow off steam "get rid of pent-up energy or emotion" During the train ride, the kids were excited to see the old locomotive blow off steam. Y
get a grip "begin to deal with or understand" He struggled to get a grip on the slippery glass jar of pickles. Y
Table 3: Examples of expert annotations. Definitions are taken from Ayto (2020). "N" and "Y" stands for "No" and "Yes", respectively.
Appendix B Additional Annotations of DICE
The analysis we focus on in this paper uses the main subsets of DICE. There are additional sentences we have collected that can be used for further analysis in the domain of idiomaticity. A summary is presented in Table 4. We make all annotations collected publicly available.
Counts Examples and Remarks All annotated sentences 2349 This includes the aforementioned 1033 literal sentences. Unique expressions 783 - Ambiguous sentences 165 The panda car is a popular item in the collectible toy market. Figurative/Idiomatic sentences 465 It was a close call when the hiker almost slipped off the cliff. Change in Form sentences 32 She reached into the bag to find her glasses. (The idiom is "in the bag".) Doesn’t make sense sentences 162 When the children play at the park, their parents always remind them to play it safe. Grammatical Error sentences 9 The old locomotive runs out of steam halfway up the mountain. Can’t be literal sentences ("skips") 462 The nurse cared for the critical patients day in, day out without a moment’s rest. Total sentences 1295 -
Table 4: Properties of the additional annotations that we have collected.
Appendix C Sentence Generation Prompt
The prompt we used for generating the sentences is shown here. For other configurations that are not mentioned, we used the default setting.
Model: GPT-4
"role": "system", "content": "You are an expert of English"
"role": "user", "content": "Generate three sentences using the expression: ’idiom’, where the expression has a literal meaning. Each sentence must contain the expression unchanged. Format these sentences as a Python list. Don’t say anything that are not the sentences."
temperature = 0.8
Appendix D Participant Briefing
Upon signing up for participation, each annotator received a 30mins training session where they were shown examples, including Table 3.
Figure 5: Parts of the briefing annotators received.
Appendix E Implementation Details
E.1 Zero-shot Prompts
We ran the FLAN-T5 models on a NVIDIA H100 GPU. We use OpenAI’s API for interactions with the GPT models, HuggingFace for Flan-T5 models and Replicate for the Llama models. Each model was evaluated with three different prompts. All of the results we report are the average across the three prompt settings.
Model: GPT-4, GPT-4o, GPT-3.5-Turbo
Prompt 1: "Is the expression ’idiom’ used figuratively or literally in the sentence: ’sentence’. Answer ’i’ for figurative, ’l’ for literal."
Prompt 2: "In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal."
Prompt 3: "How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning."
Models: Flan-T5-XXL, Flan-T5-XL, Flan-T5-Large, Flan-T5-Small
Prompt 1: "Is the meaning of expression idiomatic or literal? If used idiomatically, answer ’i’, if literally, answer ’l’." "Expression: idiom" "Sentence: sentence".
Prompt 2: "In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal."
Prompt 3: "How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning."
Models: Llama Models
Prompt 1: Is the expression ’idiom’ used figuratively or literally in the sentence: ’sentence’. Answer ’i’ for figurative, ’l’ for literal. Put your answer after ’output: ’
Prompt 2: In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal. Put your answer after ’output: ’
Prompt 3: How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning. Put your answer after ’output: ’
E.2 Fewshot Prompt
We used the same examples across all model evaluations. "oneshot_example" is appended to the end of the above detailed prompts. oneshot_example: "Here is an example: The expression ’play with fire’ is used figuratively in the sentence: ’The war took away the unfortunate necessity, as Unionists saw it, to play with fire in the national interest, but it did not materially alter their view of themselves.’ → Output: i The expression ’play with fire’ is used literally in the sentence: ’Despite the danger, he decided to play with fire, poking the embers with a stick.’ → Output: l"
E.3 Hyper-parameters
The hyper-parameters we used for running the evaluation are provided in Table 5.
Table 5: Generated by Spread-LaTeX
GPT Models Flan-T5 Models Llama 2s Llama 3s Llama 3.1
Temperature 1 - 0.7 0.7 0.6
max_tokens Default 512 512 512 512
top_p 1 - 0.95 0.95 0.9
Appendix F Additional Results for Frequency and Likelihood Analysis
We present the results we obtained across all the models. This includes both frequency analysis per setting, and across both settings.
Due to resource limitations we could not obtain probabilities for the larger models, which are Llama 3.1 (405B), Llama 3 (70B) and Llama 2 (70B). As a result, we only conducted the frequency analysis on these models, see Figure 6. Figure 9 presents side-by-side regression plots of the frequency and likelihood analysis. Similarly, Figure 8 presents the plots for the smaller Llama models. Finally, Figure 7shows the frequency analysis plots for the GPT models.
Figure 6: Left to right: Frequency analysis for Llama 3.1 (405B), Llama 3 (70B) and Llama 2 (70B).
Figure 7: Left to right: Frequency analysis for GPT-3.5 Turbo, GPT-4 and GPT-4o.
Frequency Analysis Likelihood Analysis
Llama 2 (7B) [Uncaptioned image] [Uncaptioned image]
Llama 2 (13B) [Uncaptioned image] [Uncaptioned image]
Llama 3 (8B) [Uncaptioned image] [Uncaptioned image]
Figure 8: Visualisations of the frequency and likelihood. Smaller Llama models only.
Frequency Analysis Likelihood Analysis
Flan-T5 Small [Uncaptioned image] [Uncaptioned image]
Flan-T5 Large [Uncaptioned image] [Uncaptioned image]
Flan-T5 XL [Uncaptioned image] [Uncaptioned image]
Flan-T5 XXL [Uncaptioned image] [Uncaptioned image]
Figure 9: Visualisations of the frequency and likelihood analysis. Flan-T5 models only.